# RLIE: Rule Generation with Logistic Regression, Iterative Refinement, and Evaluation for Large Language Models

### v1

@@ 5.2 How to Best Utilize Rules?
To determine the optimal strategy for leveraging the learned probabilistic rules, we evaluated the four distinct inference methods detailed in Section 3.4. One might hypothesize that an LLM, with its vast pre-trained knowledge, could synthesize the rule set, their learned weights, and even a reference prediction to produce superior judgments. However, the empirical results, presented in Table 2, are counterintuitive.
The most striking finding is that the simplest strategy, Linear-only, where the logistic regression model makes the final prediction, achieves the best performance on nearly all datasets. It outperforms all three LLM-based inference strategies in most cases, particularly in terms of F1 score where it is the top performer across all six tasks. This highlights the high quality of the rule set curated by RLIE and the effectiveness of the probabilistic combiner.
Injecting more information into the LLM does not yield consistent benefits. Providing the LLM with rule weights (LLM + Rules + Weights) over just the rule texts does not lead to stable performance gains, suggesting that the LLM struggles to effectively internalize and apply probabilistic information to refine its reasoning. Most surprisingly, providing the LLM with the linear model’s own (often correct) prediction as a reference (LLM + Rules + Weights + Linear Prediction) frequently results in performance degradation compared to simply using the linear model’s prediction directly. This suggests that the LLM, when tasked with integrating qualitative probabilities and multiple form evidences into reasoning, can be led astray, potentially overwriting a correct judgment with an incorrect one. This finding aligns with existing research indicating that LLMs can be inconsistent when required to strictly adhere to complex, explicit instructions, showing the deficiency of LLMs to perform fine controlled inference.
@@ 6 Discussion
Our hierarchical evaluation reveals a central finding: achieving stable, globally consistent reasoning by prompting an LLM with a set of explicitly weighted rules is non-trivial. The simplest “Linear-only” strategy, which uses the learned logistic regression model as the final classifier, performed best on nearly all datasets and achieved the highest F1 score across the board. In contrast, injecting the rule texts, their weights, and even the linear model’s reference prediction back into the LLM often failed to yield stable improvements and, in some cases, degraded performance. This result aligns with the observation that LLMs excel at semantic generation and interpretation but are less reliable at fine-grained, controlled probabilistic integration. This suggests that a more robust paradigm involves a clear division of labor: leveraging LLMs for local, semantic tasks such as judging individual rules and generating new candidates, while entrusting a transparent and calibratable probabilistic combiner—in our work, a logistic regression model with elastic net regularization—with the global tasks of weighting and selection.
From a broader perspective, using natural language rules as the fundamental unit of reasoning bypasses the rigid dependency on predefined predicate spaces found in traditional methods, offering greater expressiveness for open-ended semantics and unstructured data. At the same time, this approach avoids overwhelming the LLM’s context with complex probabilistic details from an entire rule set, thereby preserving its ability to make stable judgments at the single-rule level. Unlike frameworks that model interactions at the predicate level (e.g., Markov Logic Networks), we advocate for enhancing the capabilities of the global combiner while keeping the local interface—the LLM’s ternary judgments and coverage information—fixed. This linear log-odds combiner can be naturally extended to more sophisticated models, such as (i) interpretable additive models (e.g., GAMs) to capture rule interactions; (ii) message passing on a factor graph to globally harmonize rule confidences; (iii) Bayesian logistic regression with hierarchical priors (e.g., ARD, Laplace) for joint uncertainty and sparsity modeling; or (iv) techniques like Platt scaling or isotonic regression to explicitly calibrate the LLM’s ”abstain” outputs. These extensions are fully compatible with the current RLIE architecture. Overall, we propose a replicable engineering principle: LLMs are responsible for semantic generation and local judgment, while traditional probabilistic learning methods handle global aggregation and uncertainty management. On this foundation, stronger graphical or additive structures and more advanced inference algorithms can be progressively introduced to build more flexible and powerful, yet still interpretable and calibratable, neuro-symbolic systems.
@@ 7 Conclusion
In this paper, we addressed the critical challenge of learning robust and interpretable rule sets from data. While Large Language Models (LLMs) have created new possibilities for generating rules in natural language, existing methods have struggled with the principled aggregation and calibration of these rules. We introduced RLIE, a unified framework that synergizes the rule-generation capabilities of LLMs with the global optimization of probabilistic models. By integrating rule generation, regularized logistic regression, and error-driven iterative refinement, RLIE produces a compact and calibrated set of weighted, natural language rules. Our experiments demonstrate that this hybrid approach is highly effective. However, directly using LLMs to apply weighted rules degrades the performance of downstream tasks, which further unveils the limitations of applying LLMs in rule based reasoning. Our work underscores the value of a neuro-symbolic division of labor, where LLMs handle local semantic tasks and classical models manage global probabilistic reasoning. The paper investigates the potentials and limitations of using LLMs for inductive reasoning tasks, proposing a unified framework which integrates LLMs with classic probabilistic rule combination methods, paving the way for more building reliable AI and contribute to knowledge discovery.
8 Ethic Statement
This research adheres to the ICLR Code of Ethics. The datasets used in our experiments are derived from the HypoBench benchmark, which is composed of publicly available data from prior academic research. To the best of our knowledge, personally identifiable information has been anonymized in these datasets. We acknowledge that the rules generated by our method reflect the patterns and potential societal biases present in the underlying training data. While a primary goal of our work is to enhance interpretability, this also means that any learned biases will be transparent. We caution that the application of such rule-based systems, particularly in sensitive domains like mental health analysis or persuasion, should be accompanied by careful human oversight to mitigate the risk of misuse or the perpetuation of harmful stereotypes. Our work is intended to contribute to the development of more transparent and auditable AI systems.
9 Reproducibility Statement
We are committed to ensuring the reproducibility of our research. The complete source code for the RLIE framework, along with the scripts for running all experiments, will be made publicly available upon publication. All datasets used in this study are from the public HypoBench benchmark, and we will provide the exact data splits to facilitate direct comparison. A detailed description of our experimental setup, including model specifications, baseline implementations, and primary hyperparameters, is provided in Section 4. A comprehensive list of all prompts used to interact with the LLM for rule generation, judgment, and inference is included in the Appendix C. The code will be released once the paper is accepted.
@@ A.3 Dataset
Persuasive Argument Prediction (Dreddit). Given a pair of argumentative texts, the model must predict which is more persuasive. This task emphasizes linguistic features correlated with persuasiveness, such as argument structure, rhetorical devices, and emotional tone. It is particularly challenging as persuasiveness often depends on complex context and psychological factors.
@@
Appendix B LLM Usage

### v2

@@ 5.2 How to Best Utilize Rules?
Table 2: Impact of inference strategies (Accuracy / Macro-F1, $\times 10^{-2}$). E1 is the most consistently strong approach, while prompting-based variants (E2–E4) yield mixed results. The best Accuracy and Macro-F1 for each backbone and dataset are bolded, respectively.
Backbone Strategy Reviews Dreaddit Headlines Citations LLM Detect Retweets
DeepSeek V3.2 (E1) Combiner 70.9 / 70.7 82.3 / 82.3 67.0 / 67.0 64.6 / 63.0 90.7 / 90.7 65.7 / 65.6
(E2) Prompt(R) 65.7 / 63.1 77.5 / 76.6 66.9 / 66.8 60.4 / 56.2 89.6 / 89.6 64.4 / 64.3
(E3) Prompt(RW) 65.4 / 64.0 77.9 / 77.1 65.2 / 65.0 55.2 / 53.5 85.2 / 85.0 62.5 / 61.8
(E4) Prompt(RWP) 69.5 / 68.6 82.4 / 82.4 66.8 / 66.8 57.3 / 55.9 89.3 / 89.3 64.7 / 64.6
Qwen3 235B (E1) Combiner 71.5 / 71.4 79.9 / 79.9 60.6 / 60.4 61.5 / 60.6 88.3 / 88.3 66.5 / 66.5
(E2) Prompt(R) 64.5 / 64.2 77.3 / 76.7 60.0 / 59.7 60.4 / 60.3 88.4 / 88.4 64.5 / 63.2
(E3) Prompt(RW) 63.4 / 63.2 77.7 / 77.1 58.5 / 58.2 60.4 / 60.1 87.1 / 87.1 65.8 / 65.0
(E4) Prompt(RWP) 66.8 / 66.6 79.7 / 79.5 59.9 / 59.8 54.2 / 54.0 87.9 / 87.9 66.0 / 65.7
Qwen3-Next 80B (E1) Combiner 68.3 / 67.8 81.1 / 81.1 62.9 / 62.7 56.3 / 55.0 87.6 / 87.6 61.9 / 61.8
(E2) Prompt(R) 66.9 / 66.6 74.5 / 73.1 62.6 / 62.3 56.3 / 53.0 84.1 / 84.0 57.9 / 54.7
(E3) Prompt(RW) 64.7 / 63.9 76.2 / 75.2 64.1 / 64.0 55.2 / 53.2 75.9 / 74.5 57.1 / 53.0
(E4) Prompt(RWP) 67.7 / 67.4 79.5 / 79.1 62.0 / 61.9 55.2 / 53.7 87.6 / 87.6 58.9 / 56.4
To determine the optimal inference strategy for the learned probabilistic rules, we compare the four protocols (E1–E4) defined in Section 3.4 across different LLM backbones. Intuitively, an LLM might synthesize the rule texts, the learned weights, and even the combiner’s reference prediction to produce superior judgments; however, Table 2 shows that this does not yield consistent gains. Overall, E1 (Combiner) is the most stable and consistently strong strategy: it achieves the best or near-best Accuracy and Macro-F1 in most settings, generally outperforming the three prompting-based LLM inference variants. This suggests that the rule set learned by RLIE is of high quality, and that an explicit probabilistic combiner can more reliably translate multi-rule evidence into final decisions.
In contrast, injecting more information into the prompt does not reliably improve performance. Among all prompting-based variants, the rule-only method (E2) is often the strongest; adding rule weights (E3) does not lead to consistent improvements, suggesting that LLMs struggle to reliably exploit numeric or probabilistic signals under prompting. Further providing the combiner prediction as a reference (E4) is also not a dependable upgrade: while it occasionally matches or slightly exceeds E1, its performance is highly variable, rendering it unsuitable as a default strategy. Overall, when the model is tasked with jointly integrating multiple rules, weights, and reference signals, its final decision becomes sensitive to prompt structure and potential information conflicts, leading to inconsistent behavior.
Taken together, our results support a robust “division of labor”: employing LLMs primarily for proposing and refining rules, while relying on the combiner for global, weight-aware aggregation and final inference. This synergy preserves interpretability while delivering better performance.
We also assess the reliability of the confidence scores in Appendix C, and RLIE achieves significantly lower Expected Calibration Error compared to direct prompting in most cases, which confirms that RLIE effectively mitigates the overconfidence issue often observed in LLMs.
5.3 Computational Cost Analysis
We estimate the computational cost of the entire rule learning pipeline based on token usage statistics derived from our experiments (see Table 3). Compared to the robust baseline HypoGeniC, RLIE reduces token consumption by approximately 8%, demonstrating that our hard-example driven refinement is more sample-efficient than maintaining a large rule bank. Although RLIE consumes more computational resources than the lightweight IO Refinement, this additional cost is justified by our rigorous validation protocol (evaluating rules on the full validation set vs. only 10 samples in IO Refinement), which is crucial for ensuring the reliability of the learned neuro-symbolic system.
Table 3: Average computational cost per task. RLIE achieves a balance between efficiency and robustness, consuming fewer tokens than HypoGeniC while performing comprehensive validation. Energy costs are estimated by assuming a conservative energy intensity of 1.0 J/token for the MoE backbone.
Method Avg. Tokens (M) Est. Energy (kWh) Relative Cost
HypoGeniC 13.24 3.68 108%
RLIE (Ours) 12.23 3.40 100%
IO Refinement 08.45 2.35 069%
5.4 Case Study: Rule Evolution and Representation
Table 4: Iterative Rule Refinement Process (Retweet Task). RLIE refines the rule bank in an error-driven manner. At round $r=0$, we start from a randomly sampled example; at round $r>0$, we sample a hard example misclassified by the round r1r\!-\!1 model. [NEW]  newly added rules, [KEEP]  inherited active rules, and [DROP]  rules with zeroed weights. Text in bold indicates the semantic focus.
Task: Predict which of two tweets is retweeted more. Output: First vs Second (0/1).
Rd Prompt (Side-by-side) Rule Bank Update (Full Text) Metric
0 T1: … #House debates issue of #health insurance policies eliminated
T2: … #House voting on bill to restore #health insurance policies …
GT: Second (1)
[NEW] 1. Tweets that use stronger emotional language or dramatic framing …are more likely to be retweeted. (w=.035)
[NEW]
 2. Tweets that include vivid details, quotes, or specific examples outperform those that offer general summaries. (w=.008)
[DROP]
[…]
Acc: 0.625
F1: 0.625
1 T1: Andrew C. McCarthy: The Scheme behind the Obamacare Fraud …
T2: Like all swindles, Obamacare cannot work if its targeted victims …
GT: Second (1); Prev: First (0)
[NEW] 1. Tweets that prioritize clarity and immediate comprehension …are favored in news and public-affairs contexts. (w=.045)
[KEEP]
 2. Tweets that use stronger emotional language or dramatic framing …are more likely to be retweeted. (w=.034)
[DROP]
[…]
Acc: 0.680
F1: 0.679
2 T1: NEW VID!! :D ***MEET MY DAD!*** … RT ? :)
T2: what do you think about the SNEAK PEEK … ***MEET MY DAD!*** …
GT: Second (1); Prev: First (0)
[NEW] 1. Tweets that use personal voice or self-referential framing …are more retweetable when they convey authenticity. (w=.223)
[KEEP]
 2. Tweets that prioritize clarity and immediate comprehension …are favored in news contexts. (w=.178)
[DROP]
[…]
Acc: 0.700
F1: 0.699
To investigate the inner workings of RLIE, we conduct two case studies: one tracing the dynamic trajectory of rule refinement on Retweets, and another analyzing the semantic efficiency of the final rule set on Headlines.
Dynamics of Rule Evolution. We track the iterative updates of the rule set in Table 4. The process acts as a continuous error-correction loop. Initial heuristics like urgency or novelty start with non-zero weights but are effectively pruned to zero as they fail to generalize. Meanwhile, hard examples trigger the generation of more specific semantic features, such as clarity in news contexts, which are then up-weighted. This confirms that RLIE actively performs soft feature selection, evolving from coarse heuristics to robust, task-aligned discriminators.
Table 5: The rule set learned by RLIE. The logistic combiner assigns high importance to Surprise and Relatability features while effectively pruning the Urgency rule ($w=0$) via $L_{1}$ regularization.
Rule ID Wgt. Learned Rules Focus
Rule 2 0.039 Headlines that frame content as a surprising outcome… provoke curiosity. Surprise
Rule 0 0.005 Headlines that frame content as a direct, relatable scenario Relatable
Rule 1 0.000 Headlines that incorporate a sense of urgency (Pruned)
Figure 3: t-SNE of rule embeddings. RLIE learns complementary rules while HypoGeniC shows semantic noise and IO Refinement exhibits mode collapse (visualizing all refinement steps).
Efficiency of Representation. We visualize the embeddings of the final rule sets for the Headlines task using t-SNE in Figure 3 and inspect the learned rules in Table 5. HypoGeniC exhibits a scattershot distribution, retaining 20 diverse but noisy rules that yield suboptimal accuracy. IO Refinement suffers from mode collapse, where all 5 retained rules cluster tightly and redundantly focus on aggressive imagery. In contrast, RLIE achieves semantic efficiency, converging to just three widely separated rules that capture orthogonal concepts: Surprise and Relatability. Notably, the Urgency rule is correctly identified as ineffective and pruned, mirroring the trajectory seen in the Retweets case. This disentanglement allows RLIE to achieve the highest accuracy with the most compact representation.
@@ 6 Discussion
We propose a general paradigm for designing inference systems where natural-language rules serve as the basic unit of reasoning. In this setting, future improvements may stem less from injecting global probabilistic details into LLM prompting, and more from shaping a reusable interface that separates LLM-based semantics from probabilistic learning. Compared to traditional predicate-based frameworks, natural-language rules better accommodate open-ended semantics and unstructured data; however, asking an LLM to simultaneously interpret rules and perform global weighted aggregation can be sensitive to prompt phrasing and instruction competition, potentially hurting stability. A promising direction is to fix the local interface—letting the LLM provide per-rule signals (e.g., judgments or confidence scores) alongside coverage information—and to delegate global consistency, conflict resolution, and uncertainty management to a non-LLM, transparent aggregation module. This module can be further strengthened by interpretable additive models or explicit calibration techniques, enabling robust neuro-symbolic systems that combine the semantic flexibility of LLMs with the rigor of classical statistical models.
@@ 7 Conclusion
In this work, we introduced RLIE, a unified framework for learning compact, interpretable natural-language rule sets by integrating LLM-based hypothesis generation with a calibrated probabilistic combiner and error-driven refinement. Extensive experiments across diverse real-world tasks and LLM backbones demonstrate that direct inference via the learned logistic combiner yields the most robust and consistent performance, whereas re-injecting rules and weights into LLM prompting often leads to mixed results or degradation. These findings advocate for a principled neuro-symbolic division of labor: LLMs excel at local semantic operations—such as proposing hypotheses and judging applicability—while classical probabilistic models are superior for global selection, weighting, and calibration. This synergy enables the construction of reliable, auditable, and high-performance reasoning systems in natural language.
Impact Statement
This work advances machine learning by studying how LLMs can generate natural-language decision rules and how classic probabilistic modeling can combine these rules into calibrated, inspectable predictions. A key positive impact is improved transparency and controllability: explicit rule sets and learned weights make it easier to audit model behavior, attribute errors to specific rules, and iteratively refine systems to better match desired policies, which can be valuable for text classification and screening workflows where stakeholders require understandable rationales. The framework may also lower the barrier to building lightweight, modular decision systems by reusing and updating rule sets without retraining large models end-to-end. As with other data-driven approaches, risks include learning rules that encode biases or spurious correlations present in data, or over-trusting rule-based rationales in high-stakes settings. These risks can be mitigated through careful dataset curation and documentation, evaluation across subpopulations and robustness tests, and maintaining human oversight when deploying or updating rule sets.
@@ A.3 Dataset
Persuasive Argument Prediction (Dreaddit). The objective is to detect mental stress signals from Reddit posts across different communities. These posts, sourced from social media, often contain complex personal narratives that reflect an individual’s psychological state. The primary challenge lies in investigating specific linguistic features indicative of mental stress, identifying subtle patterns in the text that effectively signal the presence of distress.
@@
A.4 Hyperparameter Settings
Iterative Refinement. To balance performance and computational cost, we set the maximum number of refinement iterations $R_{max}=20$. The early stopping mechanism uses a patience of $p=3$ rounds with a minimum improvement threshold $\delta=1\times 10^{-3}$ on the validation set accuracy.
Logistic Regression. We implement the logistic combiner using the scikit-learn library. Hyperparameters for Elastic Net are selected via 5-fold stratified cross-validation on $\mathcal{S}_{train}$. We perform a grid search for the inverse regularization strength $C$ (30 logarithmically spaced values in $10^{-3},10^{1}]$) and the L1 ratio $\alpha$ (30 linearly spaced values in $[0.01,0.99]$).
LLM Generation. For the Rule Generation and Refinement stages, we prioritize stability and reproducibility. We set the generation temperature to $1\times 10^{-5}$. The sampling process uses a batch size of 20 training samples, generating 5 rules per API call, with a maximum token limit of 8000.
Appendix B Parameter Study
Table 6: Sensitivity analysis of the coverage threshold $\gamma$ on the Headline dataset (Backbone: DeepSeek-V3). Performance remains robust for $\gamma\in[0.1,0.5]$.
Coverage Threshold ($\gamma$) Accuracy F1 Score
0.1 66.7 66.7
0.2 67.0 67.0
0.3 67.3 67.2
0.4 66.5 66.5
0.5 66.5 66.5
0.6 65.5 65.5
0.7 66.5 66.5
0.8 65.6 65.6
0.9 64.3 64.2
Appendix C Calibration Analysis
To validate our claim that the logistic combiner provides calibrated probabilistic outputs, we evaluate the Expected Calibration Error (ECE) (Guo et al., 2017) of RLIE compared to the rule-injection prompting baseline (E2).
C.1 Methodology
We compute ECE using 10 equal-width bins. For both methods, ECE measures the weighted average difference between the predicted confidence and the empirical accuracy: $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})|$.
  • RLIE (Ours): The probability is directly derived from the logistic regression combiner: $p(y=1|x)=\sigma(\beta_{0}+\sum w_{j}r_{j}(x))$. Since logistic regression minimizes the negative log-likelihood, it naturally aligns predicted probabilities with empirical frequencies.
  • Prompting (E2): Since standard prompting yields binary text, we adapt it to produce probabilistic outputs for fair comparison. We constrain the LLM (DeepSeek-V3.2) to output a single token (e.g., ’A’ or ’B’) and extract the token log-probabilities. The probability of the positive class is computed via softmax: $p=\frac{\exp(\ell_{pos})}{\exp(\ell_{pos})+\exp(\ell_{neg})}$.
C.2 Results
Table 7 reports the mean ECE and standard deviation over 3 runs. RLIE achieves consistently lower ECE scores (better calibration) on 5 out of 6 datasets. Notably, on tasks like Retweets and Reviews, Prompting exhibits high calibration error (>0.45>0.45), indicating severe overconfidence (predictions are often close to 1.0 even when wrong). In contrast, RLIE maintains an ECE around 0.11, demonstrating that the learned weights serve as reliable uncertainty estimates.
Table 7: Comparison of Expected Calibration Error (ECE) on the test set (Lower is better). RLIE significantly outperforms Prompting (E2) on most tasks, providing more reliable confidence estimates.
Dataset RLIE (Ours) Prompting (E2)
ECE $\downarrow$ ECE $\downarrow$
Retweets 0.116 $\pm$ 0.041 0.480 $\pm$ 0.040
Reviews 0.113 $\pm$ 0.022 0.511 $\pm$ 0.065
Headlines 0.104 $\pm$ 0.029 0.371 $\pm$ 0.034
Dreaddit 0.226 $\pm$ 0.054 0.368 $\pm$ 0.050
Citations 0.179 $\pm$ 0.048 0.494 $\pm$ 0.261
LLM Detect 0.252 $\pm$ 0.052 0.187 $\pm$ 0.051
Appendix D LLM Usage